OpenStudy (anonymous):

Find cot θ if csc θ = square root of 5/2 and tan θ > 0.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

please! my last question :(

OpenStudy (anonymous):

@mathmale

OpenStudy (campbell_st):

for csc to be positive and tan > 0 this is a 1st quadrant angle you should know csc = 1/sin so using that property you can say \[\frac{1}{\sin} = \frac{\sqrt{5}}{2}\] now take the reciprocal of each side of the equation \[\sin(x) = \frac{2}{\sqrt{5}}\] now draw a triangle with the opposite side 2 and hypotenuse sqrt(5) so you have  find the value of a using pythagoras theorem then find tan(x) and lastly take the reciprocals to find cot(x) hope it helps

OpenStudy (anonymous):

i got 1/2, correct? @campbell_st

OpenStudy (campbell_st):

thats a good job.... well done.

OpenStudy (anonymous):

thanks(: